# Closing the Gap: Achieving Better Accuracy-Robustness Tradeoffs Against Query-Based Attacks

### 15 Dec 2023 v1

@@ Abstract
Although promising, existing defenses against query-based attacks share a common limitation: they offer increased robustness against attacks at the price of a considerable accuracy drop on clean samples. In this work, we show how to efficiently establish, at test-time, a solid tradeoff between robustness and accuracy when mitigating query-based attacks. Given that these attacks necessarily explore low-confidence regions, our insight is that activating dedicated defenses, such as RND (Qin et al., NeuRIPS 2021) and Random Image Transformations (Xie et al., ICLR 2018), only for low-confidence inputs is sufficient to prevent them. Our approach is independent of training and supported by theory. We verify the effectiveness of our approach for various existing defenses by conducting extensive experiments on CIFAR-10, CIFAR-100, and ImageNet. Our results confirm that our proposal can indeed enhance these defenses by providing better tradeoffs between robustness and accuracy when compared to state-of-the-art approaches while being completely training-free. [^1] The code is available at https://github.com/RUB-InfSec/closing˙the˙gap.
@@ Introduction
In an attempt to design defensive strategies to mitigate adversarial inputs, an arms race has been sparked in the community. A number of proposals have explored the use of randomization to improve adversarial robustness. While most randomization strategies are ineffective in the original white-box setting (Athalye et al. 2018; Athalye, Carlini, and Wagner 2018), recent findings suggest that embedding random noise within the input could effectively mitigate query-based black-box attacks (Byun, Go, and Kim 2022). As such, while randomization-based defenses could be effective in thwarting query-based attacks, they inevitably damage the classifier’s clean accuracy (Tsipras et al. 2019). For example, the random noise defense (Qin et al. 2021, NeuRIPS) (using the noise level $\sigma=0.07$ as a hyperparameter) can increase the robust accuracy against PopSkipJump (PSJA) (Simon-Gabriel, Sheikh, and Krause 2021), the strongest known black-box attack against randomized classifiers, by almost $13\%$ at the cost of a significant drop of almost $30\%$ in main task accuracy in the CIFAR-10 dataset.
@@
We show that our approach can be instantiated with existing test-time defenses without the need for re-training and show that it can strike robust tradeoffs that could not be reached otherwise with existing off-the-shelf defenses. While our approach is generic, it is naturally apt to particularly thwart decision-based attacks, as it is harder for the adversary to avoid low-confidence regions when she does not have access to the scores output by the classifier.
We conduct an extensive robustness evaluation of our approach with lightweight off-the-shelf defenses and find across-the-board improvements in robustness-accuracy tradeoffs over all considered defenses. For instance, our experiments on CIFAR-10 and CIFAR-100 show that for PSJA, one of the most powerful state-of-the-art decision-based attacks, our method improves robust accuracy by up to $8\%$ and $20\%$, with a negligible impact on clean task accuracy of at most $2\%$. For SurFree, a geometric decision-based attack, we report robustness improvements of up to $34\%$ with almost no degradation of main task accuracy.
@@
Transfer-based attacks (Dong et al. 2018; Xie et al. 2019; Naseer et al. 2021) have access to (similar) training data that has been used to train the target classifier. Using such datasets, they train a local “surrogate” model and use it to craft adversarial samples with white-box strategies against the surrogate model. Due to the transferability properties of DNNs, the generated samples often fool the original target classifier. To be effective, these attacks require knowledge of the target classifier’s architecture and/or its training data. On the other hand, query-based attacks (Brendel, Rauber, and Bethge 2018; Andriushchenko et al. 2020; Chen, Jordan, and Wainwright 2020) do not require access to the training data itself and instead interact with the target classifier to obtain predictions on inputs of their choice. By adaptively generating a sequence of images based on the classifier’s predictions, query-based attacks can derive adversarial examples with minimal distortion. Score-based attacks assume additional access to the individual probabilities of the possible classes, while decision-based attacks typically work with access to the top-1 label.
@@
Figure 1: Selected iterations (or steps) from the binary search procedure typically used in a decision-based attack. Here, a source image, classified as ’cat’, is blended with a target image, classified as ’bird’. The procedure crosses a low confidence region, before outputting a boundary sample with a slightly higher probability for the target class.
@@
The first known attack (Brendel, Rauber, and Bethge 2018) is based on a naive rejection-sampling exploration of the decision boundary to converge to a low-distortion adversarial example, which requires a significant amount of queries. Recent proposals can primarily be divided into two broader categories: gradient-based methods, such as HopSkipJump (HSJA) (Chen, Jordan, and Wainwright 2020), build a surrogate gradient inspired by zeroth-order-optimization for faster convergence, and geometry-based attacks such as SurFree (Maho, Furon, and Le Merrer 2021), that exploit geometric properties of the decision boundary and perform careful query trials along it.
@@
Existing mitigation strategies against adversarial examples can be categorized into two main groups: training- and inference-time approaches. The most popular training-time defense is adversarial training with various instantiations (Madry et al. 2019; Zhang et al. 2019), which essentially augments the training dataset with adversarial examples labeled with the semantically correct class. This process is especially costly as the adversarial images have to be generated at training time to allow their use during the training procedure. Other approaches, such as Parametric Noise Injection (PNI) (He, Rakin, and Fan 2019), inject Gaussian distributed noise upon weights in a layer-wise fashion or insert a noise layer before each convolution layer, as done in Random Self Ensemble (RSE) (Liu et al. 2018).
@@
Most proposals that aim at improving the robustness-accuracy tradeoff are focused on adversarial-training defenses based on data-augmentation methods, hence leading to high computational overhead at training time (Raghunathan et al. 2020). Other approaches allow for a “free” adjustment with a hyperparameter, without requiring any retraining after initially augmenting the model with a new model-conditional training approach (Wang et al. 2020).
@@
Given a genuine input $x_{0}\in\mathbb{R}^{d}$ predicted as $C(x_{0})=s$ (source class), $x^{\prime}$ is an adversarial sample of $x_{0}$ if it $C(x^{\prime})\neq s$ and $\lVert x^{\prime}-x_{0}\rVert_{p}<\varepsilon$ for a given distortion bound $\varepsilon\in\mathbb{R}^{+}$ and $l_{p}$ norm. Formally, we have:
@@
The objective of the adversary can be then expressed as follows:
@@
For an adversarial sample $x^{\prime}$ to be successful, it must have a low distortion and satisfy $\mathcal{A}_{x_{0}}(x^{\prime})>0$. The attacker searches for inputs $x^{\prime}$ solving the optimization problem, while she is restricted by the query budget $Q$ in his number of permitted queries to the classifier:
@@
We define a low-confidence region with respect to classifier $f$ and confidence threshold $\tau$ as the space of samples for which $f$ has confidence below $\tau$, i.e., $\{x\in X:\max_{i\in[n]}f_{i}(x)<\tau\}$.
@@
As most genuine samples are classified correctly with high confidence, the defense has minimal impact on the classification accuracy of genuine samples while still being effective at preventing attacks, as attackers have to visit the low confidence region where the defense is enabled in order to create adversarial samples.
@@
Predictions of recent model architectures are typically miscalibrated in the sense that they typically classify inputs with rather high confidence. To make our approach agnostic to the deployment instance, we calibrated our models by leveraging temperature scaling (Guo et al. 2017), i.e., using a single parameter $T>0$ to scale the logits $z$ before passing them through softmax $\sigma$ and updating the confidence prediction $p$ with $p=\sigma(\nicefrac{{z}}{{T}})$.
@@
$\min_{T}\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\textbf{1}(\hat{y}_{i}=y_{i})-\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p_{i}}\right|,$ (5)
where $n$ is the overall number of samples, $p_{i}$ is the confidence of the model when classifying image $i$ and $y_{i}$ is the corresponding label.
@@
Recall that the supervised training procedure for a generic multi-class classifier $C$ aims at finding the optimal set of parameters $\theta$ that minimize the aggregated loss over the entire training set $\mathcal{X}$: $\min_{\theta}\frac{1}{N}\sum_{i=1}^{N}L(y_{i},f(x_{i},\theta))$, for $N=|\mathcal{X}|$, sample $x\in\mathcal{X}$, ground-truth $y$, and loss-function $L$.
The optimization procedure itself can be instantiated with varying algorithms, such as SGD, which come with varying convergence guarantees. Depending on the use case, loss-function $L$ is, for example, the cross-entropy loss.
In what follows, we show that (1) genuine (i.e., non-adversarial) samples are unlikely to bear low confidence, and (2) existing query-based attacks necessarily have to investigate low-confidence regions to ensure convergence.
@@
Consider the loss function $L$ with the $n$-class cross-entropy loss:
$L(y,\mathbf{p}):=-\sum_{c=1}^{n}\mathbf{1}(y=c)\log(p_{c})$
with $\mathbf{p}$ being the probability vector output by the classifier. The indicator function $\mathbf{1}$ is defined as:
@@
$\mathrm{bd}(\mathcal{A}_{x_{0}}):=\{x\in\mathbb{R}^{d}:\mathcal{A}_{x_{0}}(x)=0\}.$ (6)
@@
Our choice of attacks is mainly motivated by a study that provides an overview of decision-based attacks and outlines a comparison of their contributions (cf. Figure 8, Appendix). We selected PSJA (Simon-Gabriel, Sheikh, and Krause 2021) as it is the only attack that considers a probabilistic classifier. It is based on gradient estimation and is a direct improvement over HSJA (Chen, Jordan, and Wainwright 2020). We further selected the SurFree attack (Maho, Furon, and Le Merrer 2021) as it is the most recent attack, which trades the costly gradient estimation step and instead exploits geometric properties of the decision boundary. We argue that our selected attacks exhibit good coverage over attacks in the literature.
@@
Our approach can be generically instantiated using any test-time defense $\mathsf{D}$. We selected three lightweight transformations to instantiate our approach: RND and RCR for the probabilistic setting and JPEG compression for the deterministic setting.
RND applies additive centered Gaussian noise to the input $f_{\mathsf{RND}}(\nu;x)=f(x+\nu r)\quad\text{for}\quad r\leftarrow\mathcal{N}(0,I),$ where the parameter $\nu\in\mathbb{R}$ controls the noise magnitude.
RCR is a function $\gamma_{\nu}$ that randomly crops and bi-linearly interpolates the image back to its original size: ${f_{\mathsf{RCR}}(\nu;x)=f(\gamma_{\nu}(x))},$ where $\nu$ denotes the cropping size.
Finally, JPEG applies the JPEG compression algorithm $\phi$ with $f_{\mathsf{JPEG}}(\nu;x)=f(\phi_{\nu}(x)),$where $\nu\in[0,100\%]$ is the quality parameter.
We include a deterministic defense to evaluate attacks, such as SurFree, that rely on deterministic classification, as they are otherwise easily defeated by probabilistic defenses.
Dataset CIFAR-10 CIFAR-100 ImageNet
$\nu$ / $\tau$ 0.5 0.6 0.7 0.8 0.9 0.97 0.99 1.0 0.5 0.6 0.7 0.8 0.9 0.97 0.99 1.0 0.3 0.4 0.5 0.6 0.7 0.9 0.97 1.0
@@
0.01 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.49 0.95 0.50 0.94 0.49 0.95 0.55 0.94 0.49 0.95 0.46 0.94 0.51 0.94 0.48 0.60 0.17 0.60 0.21 0.60 0.20 0.60 0.25 0.60 0.25 0.60 0.20 0.60 0.19 0.60 0.24 0.81 0.97 0.81 0.97 0.80 0.98 0.80 0.97 0.80 0.97 0.80 0.98 0.80 0.97 0.80 0.97
0.02 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.51 0.95 0.54 0.94 0.57 0.94 0.57 0.94 0.56 0.94 0.62 0.94 0.55 0.94 0.53 0.59 0.26 0.59 0.29 0.59 0.28 0.59 0.33 0.59 0.31 0.58 0.34 0.58 0.33 0.58 0.26 0.81 0.97 0.80 0.96 0.80 0.98 0.80 0.98 0.80 0.97 0.80 0.97 0.80 0.97 0.80 0.97
0.05 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.44 0.94 0.49 0.94 0.47 0.94 0.59 0.93 0.54 0.87 0.64 0.83 0.71 0.82 0.71 0.58 0.23 0.56 0.26 0.55 0.31 0.53 0.38 0.50 0.33 0.48 0.38 0.47 0.40 0.46 0.46 0.80 0.96 0.80 0.94 0.79 0.94 0.78 0.94 0.77 0.90 0.77 0.91 0.76 0.92 0.76 0.91
0.07 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.44 0.94 0.54 0.94 0.48 0.93 0.51 0.92 0.54 0.82 0.54 0.65 0.58 0.65 0.60 0.56 0.22 0.54 0.20 0.53 0.28 0.49 0.35 0.46 0.37 0.41 0.37 0.38 0.44 0.37 0.37 0.80 0.96 0.79 0.94 0.78 0.95 0.77 0.94 0.76 0.89 0.74 0.88 0.74 0.89 0.74 0.85
@@
29 200 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.37 0.95 0.47 0.94 0.44 0.94 0.35 0.94 0.42 0.92 0.32 0.91 0.33 0.91 0.37 0.59 0.24 0.59 0.29 0.58 0.31 0.56 0.37 0.55 0.42 0.54 0.36 0.54 0.44 0.53 0.35 0.81 0.96 0.80 0.95 0.80 0.96 0.80 0.97 0.80 0.95 0.80 0.97 0.80 0.96 0.80 0.95
27 176 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.43 0.95 0.39 0.94 0.41 0.94 0.41 0.94 0.38 0.90 0.42 0.88 0.39 0.88 0.34 0.59 0.25 0.58 0.37 0.57 0.40 0.56 0.42 0.53 0.43 0.51 0.41 0.50 0.49 0.50 0.46 0.80 0.96 0.80 0.93 0.80 0.93 0.79 0.95 0.79 0.93 0.78 0.94 0.78 0.94 0.78 0.94
26 152 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.38 0.95 0.41 0.94 0.45 0.94 0.39 0.93 0.47 0.89 0.33 0.86 0.34 0.86 0.39 0.58 0.21 0.58 0.37 0.56 0.36 0.54 0.39 0.52 0.49 0.49 0.47 0.47 0.50 0.47 0.48 0.80 0.94 0.80 0.94 0.79 0.92 0.78 0.91 0.77 0.93 0.76 0.93 0.76 0.93 0.76 0.91
25 128 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.41 0.95 0.35 0.94 0.43 0.94 0.39 0.93 0.35 0.88 0.46 0.83 0.34 0.83 0.33 0.58 0.27 0.57 0.28 0.56 0.34 0.54 0.44 0.50 0.42 0.47 0.45 0.45 0.47 0.44 0.47 0.80 0.95 0.79 0.93 0.78 0.95 0.76 0.93 0.75 0.92 0.71 0.90 0.70 0.92 0.70 0.91
22 104 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.42 0.94 0.42 0.94 0.45 0.94 0.43 0.93 0.45 0.81 0.43 0.66 0.51 0.67 0.49 0.57 0.24 0.55 0.27 0.53 0.29 0.50 0.36 0.46 0.40 0.39 0.37 0.36 0.43 0.34 0.43 0.79 0.94 0.78 0.92 0.76 0.91 0.74 0.93 0.71 0.88 0.61 0.81 0.60 0.80 0.59 0.83
18 80 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.48 0.94 0.49 0.94 0.46 0.93 0.46 0.92 0.37 0.72 0.48 0.39 0.39 0.38 0.41 0.55 0.23 0.52 0.28 0.50 0.34 0.46 0.28 0.40 0.29 0.30 0.35 0.25 0.33 0.22 0.31 0.79 0.94 0.77 0.93 0.74 0.91 0.71 0.88 0.66 0.85 0.48 0.70 0.44 0.67 0.43 0.71

### 21 Mar 2024 v2

@@ Abstract
Although promising, existing defenses against query-based attacks share a common limitation: they offer increased robustness against attacks at the price of a considerable accuracy drop on clean samples. In this work, we show how to efficiently establish, at test-time, a solid tradeoff between robustness and accuracy when mitigating query-based attacks. Given that these attacks necessarily explore low-confidence regions, our insight is that activating dedicated defenses, such as random noise defense and random image transformations, only for low-confidence inputs is sufficient to prevent them. Our approach is independent of training and supported by theory. We verify the effectiveness of our approach for various existing defenses by conducting extensive experiments on CIFAR-10, CIFAR-100, and ImageNet. Our results confirm that our proposal can indeed enhance these defenses by providing better tradeoffs between robustness and accuracy when compared to state-of-the-art approaches while being completely training-free [^1] The corresponding code is available at https://github.com/RUB-InfSec/closing˙the˙gap..
@@ Introduction
In an attempt to design defensive strategies to mitigate adversarial inputs, an arms race has been sparked in the community. A number of proposals have explored the use of randomization to improve adversarial robustness. While most randomization strategies are ineffective in the original white-box setting (Athalye et al. 2018; Athalye, Carlini, and Wagner 2018), recent findings suggest that embedding random noise within the input could effectively mitigate query-based black-box attacks (Byun, Go, and Kim 2022). As such, while randomization-based defenses could be effective in thwarting query-based attacks, they inevitably damage the classifier’s clean accuracy (Tsipras et al. 2019). For example, the random noise defense (Qin et al. 2021) (using the noise level $\sigma=0.07$ as a hyperparameter) can increase the robust accuracy against PopSkipJump (PSJA) (Simon-Gabriel, Sheikh, and Krause 2021), the strongest known black-box attack against randomized classifiers, by almost $13\%$ at the cost of a significant drop of almost $30\%$ in main task accuracy in the CIFAR-10 dataset.
@@
We show that our approach can be instantiated with existing test-time defenses without the need for retraining and show that it can strike robust tradeoffs that could not be reached otherwise with existing off-the-shelf defenses. While our approach is generic, it is naturally apt to particularly thwart decision-based attacks, as it is harder for the adversary to avoid low-confidence regions when she does not have access to the scores output by the classifier.
We conduct an extensive robustness evaluation of our approach with lightweight off-the-shelf defenses and find across-the-board improvements in accuracy-robustness tradeoffs over all considered defenses. For instance, our experiments on CIFAR-10 and CIFAR-100 show that for PSJA, one of the most powerful state-of-the-art decision-based attacks, our method improves robust accuracy by up to $9\%$ and $20\%$, with a negligible impact on clean task accuracy of at most $2\%$. For SurFree, a geometric decision-based attack, we report robustness improvements of up to $34\%$ with almost no degradation of main task accuracy.
@@
Transfer-based attacks (Dong et al. 2018; Xie et al. 2019; Naseer et al. 2021) have access to (similar) training data that has been used to train the target classifier. Using such datasets, they train a local “surrogate” model and use it to craft adversarial samples with white-box strategies against the surrogate model. Due to the transferability property of DNNs, the generated samples often fool the original target classifier. To be effective, these attacks require knowledge of the target classifier’s architecture and/or its training data. On the other hand, query-based attacks (Brendel, Rauber, and Bethge 2018; Andriushchenko et al. 2020; Chen, Jordan, and Wainwright 2020) do not require access to the training data itself and instead interact with the target classifier to obtain predictions on inputs of their choice. By adaptively generating a sequence of images based on the classifier’s predictions, query-based attacks can derive adversarial examples with minimal distortion. Score-based attacks assume additional access to the individual probabilities of the possible classes, while decision-based attacks typically work with access to the top-1 label.
@@
Figure 1: Selected iterations (or steps) from the binary search procedure typically used in a decision-based attack. Here, a source image, classified as ’cat’, is blended with a target image, classified as ’bird’. The procedure crosses a low-confidence region before outputting a boundary sample with a slightly higher probability for the target class.
@@
The first known attack (Brendel, Rauber, and Bethge 2018) is based on a naive rejection-sampling exploration of the decision boundary to converge to a low-distortion adversarial example, which requires a significant amount of queries. Recent proposals can primarily be divided into two broader categories: gradient-based methods, such as HopSkipJump (HSJA) (Chen, Jordan, and Wainwright 2020), build a surrogate gradient inspired by zeroth-order-optimization for faster convergence, and geometry-based attacks, such as SurFree (Maho, Furon, and Le Merrer 2021), exploit geometric properties of the decision boundary and perform careful query trials along it.
@@
Existing mitigation strategies against adversarial examples can be categorized into two main groups: training- and inference-time approaches. The most popular training-time defense is adversarial training with various instantiations (Madry et al. 2018; Zhang et al. 2019), which essentially augments the training dataset with adversarial examples labeled with the semantically correct class. This process is especially costly as the adversarial images have to be generated at training time to allow their use during the training procedure. Other approaches, such as Parametric Noise Injection (PNI) (He, Rakin, and Fan 2019), inject Gaussian distributed noise upon weights in a layer-wise fashion or insert a noise layer before each convolution layer, as done in Random Self Ensemble (RSE) (Liu et al. 2018).
@@
Most proposals that aim at improving the accuracy-robustness tradeoff are focused on adversarial-training defenses based on data-augmentation methods, hence leading to high computational overhead at training time (Raghunathan et al. 2020). Other approaches allow for a “free” adjustment with a hyperparameter, without requiring any retraining after initially augmenting the model with a new model-conditional training approach (Wang et al. 2020).
@@
Given a genuine input $x_{0}\in\mathbb{R}^{d}$ predicted as ${C(x_{0})=s}$ (source class), $x^{\prime}$ is an adversarial sample of $x_{0}$ if $C(x^{\prime})~{}\neq~{}s$ and $\lVert x^{\prime}-x_{0}\rVert_{p}\leq\varepsilon$ for a given distortion bound $\varepsilon\in\mathbb{R}^{+}$ and $l_{p}$ norm. Formally, we have:
@@
The objective of the adversary can then be expressed as follows:
@@
For an adversarial sample $x^{\prime}$ to be successful, it must have a low distortion and satisfy $\mathcal{A}_{x_{0}}(x^{\prime})>0$. The attacker searches for inputs $x^{\prime}$ solving the optimization problem, while she is restricted by the query budget $Q$ in her number of permitted queries to the classifier:
@@
We define a low-confidence region with respect to classifier $f$ and confidence threshold $\tau$ as the space of samples for which $f$ has confidence below $\tau$, i.e., $\{x\in\mathcal{X}:\max_{i\in[n]}f_{i}(x)<\tau\}$.
@@
As most genuine samples are classified correctly with high confidence, the defense has minimal impact on the classification accuracy of genuine samples while still being effective at preventing attacks, as attackers have to visit the low-confidence region where the defense is enabled in order to create adversarial samples.
@@
Predictions of recent model architectures are typically miscalibrated in the sense that they usually classify inputs with rather high confidence. To make our approach agnostic to the deployment instance, we calibrated our models by leveraging temperature scaling (Guo et al. 2017), i.e., using a single parameter $T>0$ to scale the logits $z$ before passing them through softmax $\sigma$ and updating the confidence prediction $p$ with $p=\sigma(\nicefrac{{z}}{{T}})$.
@@
$\min_{T}\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\textbf{1}(\hat{y}_{i}=y_{i})-\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p_{i}}\right|$ (5)
where $n$ is the overall number of samples, $y_{i}$ the true class label, $\hat{y}_{i}$ the predicted class label, and $\hat{p}_{i}$ the confidence of the model when classifying image $i$. The indicator function $\mathbf{1}$ is defined as:
@@
Recall that the supervised training procedure for a generic multi-class classifier $C$ aims at finding the optimal set of parameters $\theta$ that minimize the aggregated loss over the entire training set $\mathcal{X}$: $\min_{\theta}\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(y_{i},f(x_{i},\theta))$, for $N=|\mathcal{X}|$, sample $x$, ground-truth $y$, and loss-function $\mathcal{L}$.
The optimization procedure itself can be instantiated with varying algorithms, such as SGD, which come with varying convergence guarantees. Depending on the use case, loss-function $\mathcal{L}$ is, for example, the cross-entropy loss.
In what follows, we show that (1) genuine (i.e., non-adversarial) samples are unlikely to bear low-confidence, and (2) existing query-based attacks necessarily have to investigate low-confidence regions to ensure convergence.
@@
Consider the loss function $\mathcal{L}$ with the $n$-class cross-entropy loss:
$\mathcal{L}(y,\mathbf{p}):=-\sum_{c=1}^{n}\mathbf{1}(y=c)\log(p_{c})$ (6)
with $\mathbf{p}$ being the probability vector output by the classifier.
@@
$\mathrm{bd}(\mathcal{A}_{x_{0}}):=\{x\in\mathbb{R}^{d}:\mathcal{A}_{x_{0}}(x)=0\}$ (7)
@@
Our choice of attacks is mainly motivated by a study that provides an overview of decision-based attacks and outlines a comparison of their contributions. The study can be found in the Appendix. We selected PSJA (Simon-Gabriel, Sheikh, and Krause 2021) as it is the only attack that considers a probabilistic classifier. It is based on gradient estimation and is a direct improvement over HSJA (Chen, Jordan, and Wainwright 2020). We further selected the SurFree attack (Maho, Furon, and Le Merrer 2021) as it is the most recent attack, which trades the costly gradient estimation step and instead exploits geometric properties of the decision boundary. We argue that our selected attacks exhibit good coverage over attacks in the literature.
@@
Our approach can be generically instantiated using any inference-time defense $\mathsf{D}$. We selected three lightweight transformations to instantiate our approach: RND and RCR for the probabilistic setting and JPEG compression for the deterministic setting.
RND applies additive centered Gaussian noise to the input $f_{\mathsf{RND}}(\nu,x)=f(x+\nu r)\quad\text{for}\quad r\leftarrow\mathcal{N}(0,I),$ where the parameter $\nu\in\mathbb{R}$ controls the noise magnitude.
RCR is a function $\gamma_{\nu}$ that randomly crops and bi-linearly interpolates the image back to its original size: ${f_{\mathsf{RCR}}(\nu,x)=f(\gamma_{\nu}(x))},$ where $\nu$ denotes the cropping size.
Finally, JPEG applies the JPEG compression algorithm $\phi$ with $f_{\mathsf{JPEG}}(\nu,x)=f(\phi_{\nu}(x)),$where $\nu\in[0,100\%]$ is the quality parameter. We include a deterministic defense to evaluate attacks, such as SurFree, that rely on deterministic classification, as they are otherwise easily defeated by probabilistic defenses (Chen, Jordan, and Wainwright 2020).
Dataset CIFAR-10 CIFAR-100
$\nu$ / $\tau$ 0.5 0.6 0.7 0.8 0.97 0.99 1.0 0.5 0.6 0.7 0.8 0.9 0.97 0.99 1.0
@@
0.02 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.51 0.95 0.54 0.94 0.57 0.94 0.57 0.94 0.62 0.94 0.55 0.94 0.53 0.59 0.26 0.59 0.29 0.59 0.28 0.59 0.33 0.59 0.31 0.58 0.34 0.58 0.33 0.58 0.26
0.05 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.44 0.94 0.49 0.94 0.47 0.94 0.59 0.87 0.64 0.83 0.71 0.82 0.71 0.58 0.23 0.56 0.26 0.55 0.31 0.53 0.38 0.50 0.33 0.48 0.38 0.47 0.40 0.46 0.46
0.07 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.44 0.94 0.54 0.94 0.48 0.93 0.51 0.82 0.54 0.65 0.58 0.65 0.60 0.56 0.22 0.54 0.20 0.53 0.28 0.49 0.35 0.46 0.37 0.41 0.37 0.38 0.44 0.37 0.37
@@
29 200 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.37 0.95 0.47 0.94 0.44 0.94 0.35 0.92 0.32 0.91 0.33 0.91 0.37 0.59 0.24 0.59 0.29 0.58 0.31 0.56 0.37 0.55 0.42 0.54 0.36 0.54 0.44 0.53 0.35
27 176 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.43 0.95 0.39 0.94 0.41 0.94 0.41 0.90 0.42 0.88 0.39 0.88 0.34 0.59 0.25 0.58 0.37 0.57 0.40 0.56 0.42 0.53 0.43 0.51 0.41 0.50 0.49 0.50 0.46
26 152 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.38 0.95 0.41 0.94 0.45 0.94 0.39 0.89 0.33 0.86 0.34 0.86 0.39 0.58 0.21 0.58 0.37 0.56 0.36 0.54 0.39 0.52 0.49 0.49 0.47 0.47 0.50 0.47 0.48
25 128 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.41 0.95 0.35 0.94 0.43 0.94 0.39 0.88 0.46 0.83 0.34 0.83 0.33 0.58 0.27 0.57 0.28 0.56 0.34 0.54 0.44 0.50 0.42 0.47 0.45 0.45 0.47 0.44 0.47
22 104 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.42 0.94 0.42 0.94 0.45 0.94 0.43 0.81 0.43 0.66 0.51 0.67 0.49 0.57 0.24 0.55 0.27 0.53 0.29 0.50 0.36 0.46 0.40 0.39 0.37 0.36 0.43 0.34 0.43
18 80 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.48 0.94 0.49 0.94 0.46 0.93 0.46 0.72 0.48 0.39 0.39 0.38 0.41 0.55 0.23 0.52 0.28 0.50 0.34 0.46 0.28 0.40 0.29 0.30 0.35 0.25 0.33 0.22 0.31